2. Valle de Estós

’BALL DE LLITEROLA’ (LLITEROLA VALLEY) Recommended to spend half day (morning).

The departure point is the Baños de Benasque (Thermal Baths). Cross the main road river and you will find a silverfir wood. After 25 minutes you will get to the Plleta del Llosero. Follow the open track on the right of the valley. After a hard steep incline, keep going to the bottom of the valley until you get to the refuge and the Plleta de Lliterola after 1 hour. The altitude there is 2,015 m.

From that point, if you follow the ravine to the right, in 1 hour you will get to the Ibonet de Lliterola (2,460 m.) where you will find a very inclined slope that will take you to the Ibón Blanco (2,730 m. high).